OpenStudy (anonymous):

Find the inverse laplace transform of e^(-10s)/(s(s^2+3s+2)).

OpenStudy (loser66):

hey, for this problem partial fraction decompose is what I shud say, hehe

OpenStudy (anonymous):

So it's e^(-10s)/s+e^(-10s)/(s^2+3s+2)?

OpenStudy (loser66):

the last term can be fraction again, girl

OpenStudy (anonymous):

find the inverse of: \[\frac{ 1 }{ s(s+1)(s+2) }\] the e^(as) makes you modify each term to F(s-a)

OpenStudy (anonymous):

So you do A/s+B/(s+1)+C/(s+2) and you find A, B, C?

OpenStudy (anonymous):

yes

OpenStudy (loser66):

don't forget time e^(-10s) to each term

OpenStudy (anonymous):

Wait a minute.

OpenStudy (loser66):

nope, hehehe

OpenStudy (anonymous):

Mean woman...

OpenStudy (anonymous):

sorry made a mistake

OpenStudy (loser66):

hahaha....

OpenStudy (anonymous):

the e^(as) ALSO multiplies each terms by: \[U(t - a)\] where U is the unit step function

OpenStudy (anonymous):

and i meant turn the inverse into f(t-a)

OpenStudy (anonymous):

after finding f(t) of everything except e^(as) in F(s)

OpenStudy (anonymous):

But how do I find A, B, C?

OpenStudy (anonymous):

\[\frac{ 1 }{ s(s+1)(s+2) } = \frac{ A }{ s} + \frac{ B }{ s+1 } + \frac{ C }{ s+2 }\] The partial fraction decomposition is another way of writing the fraction. If you've ever added fractions, you know that they need to have the same denominator. so make them all the same. \[\frac{ 1 }{ s(s+1)(s+2) } = \frac{ A(s+1)(s+2) }{ s(s+1)(s+2) } + \frac{ B(s)(s+2) }{s(s+1)(s+2) } + \frac{ C(s)(s+1) }{ s(s+1)(s+2) }\] \[\frac{ 1 }{ s(s+1)(s+2) } = \frac{ expand }{ s(s+1)(s+2) }\]

OpenStudy (loser66):

yes, thank a lot Euler271

OpenStudy (anonymous):

and we are saying that 1 is equal to the expanded numerator which is in terms of A, B and C

OpenStudy (anonymous):

let me know